Let a and b be the roots of the equation x^2+x+1=0. The equation whose roots are a^19, b^7 is. A. x^2-x-1=0. B. x^2-x+1=0. C. x^2+x-1=0. D. x^2+x+1=0.

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
ab=1, a+b=-1
a=1e^(2pi/3)
b=1e^(-2pi/3)
a^19b^7=1
a^19+b^7=-1
=>x^2+x+1
from above analysis , x^2+x+1 is the correct answer
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty